OpenStudy (anonymous):

if x=t^3-t and y = dqrt(3t+1), then dy/dx at t = 1 is

OpenStudy (anonymous):

dy/dx = (dy/dt)/(dx/dt)

OpenStudy (freckles):

\[\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}\]

OpenStudy (freckles):

lol you beat me

OpenStudy (anonymous):

:P coz i wasn't type laTex